a cyclist traveled at a rate of 32mph to visit a near by town. The cyclist averaged 10mph on the return trip. If the round trip took 6.3 hours, find the distance to the nearby town.
D=6.3/(32-10)
D=.286363
This is not correct, but I think the opportunity is in the set up.
Thanks for helping in advance.

Answer by checkley77(12569)

D=RT
SEEING AS THE DISTANCES ARE THE SAME WE SET THE TWO DISTANCES EQUAL TO EACH OTHER.
32X=10(6.3-X)
32X=63-10X
32X+10X=63
42X=63
X=63/42
X=1.5 HOURS FOR THE 32 MPH TRIP.
6.3-1.5=4.8 HOURS FOR THE RETURN TRIP.
32*1.5=10*4.8
48=48 MILES IS THE DISTANCE TO THE NEARBY TOWN.